OpenStudy (anonymous):

What are the next two terms in the sequence: 7,2,5,-3,8,-11,_,_

OpenStudy (campbell_st):

its similar to a fibonacci sequence. It alternates positive to negative and uses the previous 2 terms. 7 - 5 = 2 2 + 3 = 5 5 -( 5 + 3) = -3 -3 + (8 + 3) = 8 8 - (8 + 11) = -11 -11 + ( 11 + 19) = 19 19 - ( 19 + 30) = - 30

OpenStudy (anonymous):

thank you